# Improving governance outcomes through AI documentation: Bridging theory and practice

### v1

@@ Acknowledgements.
- Ahlawat et al. (2024) Archana Ahlawat, Amy Winecoff, and Jonathan Mayer. 2024. Minimum Viable Ethics: From Institutionalizing Industry AI Governance to Product Impact. arXiv (September 11 2024). https://www.arxiv.org/abs/2409.06926
@@
- Blasch et al. (2021) Erik Blasch, James Sung, and Tao Nguyen. 2021. Multisource AI scorecard table for system evaluation. In AAAI FSS20: Artificial Intelligence in Government and Public Sector.
@@
- Chang and Custis (2022) Jiyoo Chang and Christine Custis. 2022. Understanding Implementation Challenges in Machine Learning Documentation. In Equity and Access in Algorithms, Mechanisms, and Optimization. ACM, Arlington VA USA, 1–8. https://doi.org/10.1145/3551624.3555301
@@
- Chmielinski et al. (2020) Kasia S Chmielinski, Sarah Newman, Matt Taylor, Josh Joseph, Kemi Thomas, Jessica Yurkofsky, and Yue Chelsea Qiu. 2020. The Dataset Nutrition Label (2nd Gen): Leveraging Context to Mitigate Harms in Artificial Intelligence.
@@
- Geiger et al. (2020) R Stuart Geiger, Kevin Yu, Yanlai Yang, Mindy Dai, Jie Qiu, Rebekah Tang, and Jenny Huang. 2020. Garbage in, garbage out? Do machine learning application papers in social computing report where human-labeled training data comes from?. In Proceedings of the 2020 conference on fairness, accountability, and transparency. 325–336.
@@
- Green et al. (2021) Nekesha Green, Chavez Procope, Adeel Cheema, and Adekunle Adediji. 2021. System Cards, a New Resource for Understanding How AI Systems Work. https://ai.facebook.com/blog/system-cards-a-new-resource-for-understanding-how-ai-systems-work/
@@
- Holland et al. (2018) Sarah Holland, Ahmed Hosny, Sarah Newman, Joshua Joseph, and Kasia Chmielinski. 2018. The Dataset Nutrition Label: A Framework To Drive Higher Data Quality Standards.
@@
- Liao and Vaughan (2023) Q Vera Liao and Jennifer Wortman Vaughan. 2023. Ai transparency in the age of llms: A human-centered research roadmap. arXiv preprint arXiv:2306.01941 (2023), 5368–5393.
@@
- Palinkas et al. (2015) Lawrence A Palinkas, Sarah M Horwitz, Carla A Green, Jennifer P Wisdom, Naihua Duan, and Kimberly Hoagwood. 2015. Purposeful sampling for qualitative data collection and analysis in mixed method implementation research. Administration and policy in mental health and mental health services research 42 (2015), 533–544.
@@
- Pepe et al. (2024) Federica Pepe, Vittoria Nardone, Antonio Mastropaolo, Gerardo Canfora, and Gabriele Bavota. 2024. How Do Hugging Face Models Document Datasets, Bias, and Licenses? An Empirical Study. (2024).
@@
- Procope et al. (2022) Chavez Procope, Adeel Cheema, David Adkins, Bilal Alsallakh, Emily McReynolds, Grace Pehl, Erin Wang, and Polina Zvyagina. 2022. System-Level Transparency of Machine Learning. (February 22 2022). https://ai.meta.com/research/publications/system-level-transparency-of-machine-learning/
@@
- Srinivasan et al. (2021) Ramya Srinivasan, Emily Denton, Jordan Famularo, Negar Rostamzadeh, Fernando Diaz, and Beth Coleman. 2021. Artsheets for Art Datasets.

### v2

@@ Acknowledgements.
- Ahlawat et al. (2024) Archana Ahlawat, Amy Winecoff, and Jonathan Mayer. 2024. Minimum Viable Ethics: From Institutionalizing Industry AI Governance to Product Impact. arXiv preprint arXiv:2409.06926 (2024). https://arxiv.org/html/2409.06926v1
@@
- Balayn et al. (2024) Agathe Balayn, Lorenzo Corti, Fanny Rancourt, Fabio Casati, and Ujwal Gadiraju. 2024. Understanding Stakeholders’ Perceptions and Needs Across the LLM Supply Chain. arXiv preprint arXiv:2405.16311 (2024). https://arxiv.org/abs/2405.16311
- Balayn et al. (2023) Agathe Balayn, Mireia Yurrita, Jie Yang, and Ujwal Gadiraju. 2023. ”Fairness Toolkits, A Checkbox Culture?” On the Factors that Fragment Developer Practices in Handling Algorithmic Harms. In Proceedings of the 2023 AAAI/ACM Conference on AI, Ethics, and Society. 482–495. https://dl.acm.org/doi/10.1145/3600211.3604674
@@
- Bhatt et al. (2020) Umang Bhatt, Alice Xiang, Shubham Sharma, Adrian Weller, Ankur Taly, Yunhan Jia, Joydeep Ghosh, Ruchir Puri, José MF Moura, and Peter Eckersley. 2020. Explainable Machine Learning in Deployment. In Proceedings of the 2020 Conference on Fairness, Accountability, and Transparency. 648–657. https://arxiv.org/abs/1909.06342
- Blasch et al. (2021) Erik Blasch, James Sung, and Tao Nguyen. 2021. Multisource AI Scorecard Table for System Evaluation. In AAAI FSS20: Artificial Intelligence in Government and Public Sector. https://arxiv.org/abs/2102.03985
- Bommasani et al. (2023) Rishi Bommasani, Kevin Klyman, Shayne Longpre, Sayash Kapoor, Nestor Maslej, Betty Xiong, Daniel Zhang, and Percy Liang. 2023. The foundation model transparency index. arXiv preprint arXiv:2310.12941 (2023). https://arxiv.org/abs/2310.12941
@@
- Chang and Custis (2022) Jiyoo Chang and Christine Custis. 2022. Understanding Implementation Challenges in Machine Learning Documentation. In Proceedings of the 2nd ACM Conference on Equity and Access in Algorithms, Mechanisms, and Optimization. ACM, Arlington VA USA, 1–8. https://doi.org/10.1145/3551624.3555301
@@
- Chien et al. (2023) Andrew A Chien, Liuzixuan Lin, Hai Nguyen, Varsha Rao, Tristan Sharma, and Rajini Wijayawardana. 2023. Reducing the Carbon Impact of Generative AI Inference (today and in 2035). In Proceedings of the 2nd Workshop on Sustainable Computer Systems. 1–7. https://dl.acm.org/doi/10.1145/3604930.3605705
@@
- Chmielinski et al. (2020) Kasia S Chmielinski, Sarah Newman, Matt Taylor, Josh Joseph, Kemi Thomas, Jessica Yurkofsky, and Yue Chelsea Qiu. 2020. The Dataset Nutrition Label (2nd Gen): Leveraging Context to Mitigate Harms in Artificial Intelligence. https://arxiv.org/abs/2201.03954
@@
- Dotan et al. (2024) Ravit Dotan, Borhane Blili-Hamelin, Ravi Madhavan, Jeanna Matthews, and Joshua Scarpino. 2024. Evolving AI Risk Management: A Maturity Model based on the NIST AI Risk Management Framework. arXiv preprint arXiv:2401.15229 (2024). https://arxiv.org/abs/2401.15229
@@
- Geertz (2008) Clifford Geertz. 2008. Thick Description: Toward an Interpretive Theory of Culture. In The Cultural Geography Reader. Routledge, 41–51.
- Geiger et al. (2020) R Stuart Geiger, Kevin Yu, Yanlai Yang, Mindy Dai, Jie Qiu, Rebekah Tang, and Jenny Huang. 2020. Garbage in, Garbage out? Do Machine Learning Application Papers in Social Computing Report Where Human-labeled Training Data Comes From?. In Proceedings of the 2020 Conference on Fairness, Accountability, and Transparency. 325–336. https://arxiv.org/abs/1912.08320
@@
- Glaser (1978) Barney G Glaser. 1978. Theoretical sensitivity. University of California,.
- Glaser et al. (1968) Barney G Glaser, Anselm L Strauss, and Elizabeth Strutzel. 1968. The Discovery of Grounded Theory: Strategies for Qualitative Research. Nursing Research 17, 4 (1968), 364.
@@
- Holland et al. (2018) Sarah Holland, Ahmed Hosny, Sarah Newman, Joshua Joseph, and Kasia Chmielinski. 2018. The Dataset Nutrition Label: A Framework To Drive Higher Data Quality Standards. https://arxiv.org/abs/1805.03677
@@
- Liao and Vaughan (2024) Q Vera Liao and Jennifer Wortman Vaughan. 2024. AI Transparency in the Age of LLMs: A Human-Centered Research Roadmap. Harvard Data Science Review (2024), 1–53. https://hdsr.mitpress.mit.edu/pub/aelql9qy/release/2
@@
- Palinkas et al. (2015) Lawrence A Palinkas, Sarah M Horwitz, Carla A Green, Jennifer P Wisdom, Naihua Duan, and Kimberly Hoagwood. 2015. Purposeful Sampling for Qualitative Data Collection and Analysis in Mixed Method Implementation Research. Administration and Policy in Mental Health and Mental Health Services Research 42 (2015), 533–544. https://pmc.ncbi.nlm.nih.gov/articles/PMC4012002/
@@
- Pepe et al. (2024) Federica Pepe, Vittoria Nardone, Antonio Mastropaolo, Gerardo Canfora, and Gabriele Bavota. 2024. How Do Hugging Face Models Document Datasets, Bias, and Licenses? An Empirical Study. Proceedings of the 32nd IEEE/ACM International Conference on Program Comprehension (2024). https://mdipenta.github.io/files/icpc2024.pdf
@@
- Procope et al. (2022) Chavez Procope, Adeel Cheema, David Adkins, Bilal Alsallakh, Emily McReynolds, Grace Pehl, Erin Wang, and Polina Zvyagina. 2022. System-Level Transparency of Machine Learning. https://ai.meta.com/research/publications/system-level-transparency-of-machine-learning/
@@
- Qi et al. (2023) Xiangyu Qi, Yi Zeng, Tinghao Xie, Pin-Yu Chen, Ruoxi Jia, Prateek Mittal, and Peter Henderson. 2023. Fine-Tuning Aligned Language Models Compromises Safety, Even When Users Do Not Intend To! arXiv (October 5 2023). http://arxiv.org/abs/2310.03693
@@
- Rakova et al. (2021) Bogdana Rakova, Jingying Yang, Henriette Cramer, and Rumman Chowdhury. 2021. Where responsible AI meets reality: Practitioner perspectives on enablers for shifting organizational practices. Proceedings of the ACM on Human-Computer Interaction 5, CSCW1 (2021), 1–23. https://dl.acm.org/doi/10.1145/3449081
@@
- Sambasivan et al. (2021) Nithya Sambasivan, Shivani Kapania, Hannah Highfill, Diana Akrong, Praveen Paritosh, and Lora M Aroyo. 2021. “Everyone wants to do the model work, not the data work”: Data Cascades in High-Stakes AI. In proceedings of the 2021 CHI Conference on Human Factors in Computing Systems. 1–15. https://dl.acm.org/doi/abs/10.1145/3411764.3445518
@@
- Solaiman et al. (2023) Irene Solaiman, Zeerak Talat, William Agnew, Lama Ahmad, Dylan Baker, Su Lin Blodgett, Canyu Chen, Hal Daumé III, Jesse Dodge, Isabella Duan, et al. 2023. Evaluating the Social Impact of Generative AI Systems in Systems and Society. arXiv preprint arXiv:2306.05949 (2023). https://arxiv.org/abs/2306.05949
- Srinivasan et al. (2021) Ramya Srinivasan, Emily Denton, Jordan Famularo, Negar Rostamzadeh, Fernando Diaz, and Beth Coleman. 2021. Artsheets for Art Datasets. In Proceedings of the Neural Information Processing Systems Track on Datasets and Benchmarks 1 (NeurIPS Datasets and Benchmarks 2021). https://datasets-benchmarks-proceedings.neurips.cc/paper/2021/hash/9b8619251a19057cff70779273e95aa6-Abstract-round2.html
@@
- Suresh et al. (2021) Harini Suresh, Steven R Gomez, Kevin K Nam, and Arvind Satyanarayan. 2021. Beyond expertise and roles: A Framework to Characterize the Stakeholders of Interpretable Machine Learning and Their Needs. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems. 1–16. https://dl.acm.org/doi/10.1145/3411764.3445088
- Suri (2011) Harsh Suri. 2011. Purposeful sampling in qualitative research synthesis. Qualitative research journal 11, 2 (2011), 63–75.
@@
- Weidinger et al. (2023) Laura Weidinger, Maribeth Rauh, Nahema Marchal, Arianna Manzini, Lisa Anne Hendricks, Juan Mateos-Garcia, Stevie Bergman, Jackie Kay, Conor Griffin, Ben Bariach, et al. 2023. Sociotechnical Safety Evaluation of Generative AI Systems. arXiv preprint arXiv:2310.11986 (2023). https://arxiv.org/abs/2310.11986
@@
- Yurrita et al. (2023) Mireia Yurrita, Agathe Balayn, and Ujwal Gadiraju. 2023. Generating Process-Centric Explanations to Enable Contestability in Algorithmic Decision-Making: Challenges and Opportunities. arXiv preprint arXiv:2305.00739 (2023). https://arxiv.org/abs/2305.00739